OpenStudy (lucaz):

find the solution set x^2e^x = 2xe^x

OpenStudy (zehanz):

This is of the form \(a\cdot b=a \cdot c\). The solution of this form is \(a=0 \vee b=c\). You have: \(x^2e^x=2xe^x\), so you could rewrite it as (because \(a=e^x,~ b=x^2,~ c=2x\)): \(e^x=0\vee x^2=2x \) So can you do the next step?

OpenStudy (lucaz):

I'm a bit confused. exponential equations both sides must have the same base right?

OpenStudy (zehanz):

Point is, it is not only an exponential equation, there are other factors in the equation. What I explained is, you can split up into two separate, simpler equations: \(e^x=0\) \(x^2=2x\)

OpenStudy (zehanz):

Now, because \(e^x >0\), for all values of x, we can forget this part! So we only have to solve \(x^2=2x\), or \(x^2-2x=0\)

OpenStudy (lucaz):

ok, but what this "V" means?

OpenStudy (zehanz):

Sorry, it just means "or" :)

OpenStudy (lucaz):

oh, right. thanks

OpenStudy (zehanz):

YW! Can you find the solution(s) of \(x^2-2x=0\)?

OpenStudy (lucaz):

x(x-2)=0 {0,2} ?

OpenStudy (zehanz):

Yup!

OpenStudy (lucaz):

ok, thank you

OpenStudy (zehanz):

Welcome!